Circular No. 7581

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2001 CV_{8} (LINEAR) An object that was reported as asteroidal by LINEAR (discovery obser-

vation given below), and given the designation 2001 CV_{8}, has been found by other CCD observers to show cometary activity. M. Hicks, Jet Propulsion Laboratory, reports that nonphotometric images (with imperfect tracking) obtained with the 0.61-m f/16 reflector at Table Mountain Observatory by D. Esqueda, A. Esqueda, and T. H. Ha on Feb. 4 indicate this object to be diffuse without condensation but with a faint, 5′′ fan-shaped tail toward the west. Images taken by D. T. Durig (Sewanee, TN; 0.3-m f/5.8 Schmidt-Cassegrain telescope; moonlight and tracking problems) on Feb. 6 show the object to be more diffuse than nearby stars. Observations by J. Tichá and M. Tichý at Kleť (0.57-m f/5.2 reflector) on Feb. 10 show the object to be diffuse with a 9′′ coma and a faint 15′′ tail in p.a. 270^{o}.

2001 UT  α_{2000}  δ_{2000}  m_{2}
Feb. 1.34987  10^{h}58^{m}18.^{s}71  +15^{o}47′43′′.3  19.0

Additional astrometry and the following orbital elements (and an ephemeris) by B. G. Marsden appear on MPEC 2001-C24. The elements indicate that the comet passed ∼ 0.14 AU from Jupiter in Nov. 1998.

T = 2001 Jan. 23.3533 TT  ω = 142.^{o}8087
e = 0.459019  Ω = 0.8554 }2000.0
q = 2.122429 AU  i = 8.6908
a = 3.923299 AU n^{o} = 0.1268315 P = 7.771 years

SUPERNOVA 2001T IN MCG –02-37-6 X. J. Jiang, Y. L. Qiu, and J. Y. Hu, on behalf of Beijing Astronomical

Observatory (BAO) Supernova Survey, report the discovery of an apparent supernova found on unfiltered images taken with the BAO 0.6-m telescope on Feb. 6.9 (mag ∼ 17.0) and 8.9 UT (mag ∼ 16.8). SN 2001T is located at α = 14^{h}32^{m}53^{s}.17, δ = −12^{o}58′43′′.6 (equinox 2000.0), which is 3′′.1 east and 12′′.6 south of the nucleus of MCG –02-37-6. An unfiltered image of the same field taken on 2000 June 4.7 showed nothing at this position (limiting mag ∼ 18.5).

2001 February 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green